

********************************************************************
UARS WEEKLY STATUS REPORT - 13 February 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems.
***********
PEM
***********
PEM is currently off.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
HALOE has been off during the present graze period. HALOE is scheduled to
turn on midday February 12, 1998.
***********
MLS
***********
MLS was reactivated successfully on Jan 30, in the reduced 85 watt power
mode. MLS instrument operations have continued flawlessly. MLS was powered
off at 043:12:27 UT. MLS is scheduled to resume north viewing operations
starting Feb 20.
***********
HRDI
***********
HRDI is currently on and operating normally.
***********
WINDII
***********
The WINDII Team is currently assessing the results of the production
runs with the Version 5 software.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
Testing and configuration of the DEC ALPHA workstation, which will
replace the VAX-4000 system, is in progress.
***********
CDHF
***********
The CDHF has received a total of 368 CD-ROM requests as of 13-February-1998.
There was one order received and one order shipped.
The CDHF previewed current day DAAC data for HALOE, HRDI, PEM, NMC, AND UKMO
for the month of December 1997.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-904C=475L)
119 days MLS (4.90) PROD (1344R-442C-3F=902L)
0 day CLAES (7.0) ENG (17R-1F=17L)
1 day MLS (4.22) PROD (1R-1C=0L)
CDHF pending software upgrades: IDL 5.0, OpenIngres, UCSS.
***********
FOT
***********
REPORTING PERIOD
This report covers 31 January 1998 (Orbit 34911, GMT Day 031) through 06
February 1998 (Orbit 35016, GMT Day 037).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 032/2330 34941 AXIS off
PEM 033/2231 34955 ZEPS off
WINDII 034/0009 34956 On
WINDII 034/0808 34961 Off
WINDII 034/1324 34964 On
The beta angle reached a maximum angle of 72.6 degrees on 05 February 1998
(DOY 036). The beta angle is now decreasing toward a minimum angle of 0.0
degrees on 25 February 1998 (DOY 056).
Six SOLSTICE offsets requested by SUSIM operations were performed this
week. One offset took place each day, beginning with GMT Day 031 and
ending with GMT Day 036.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 26.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta between
Battery 2 and Battery 3 of 1.3 degrees C. Additional battery performance
data for each battery on 31 January 1998 and 06 February 1998 are provided
in Appendix D.
The clock error ranged from -9.2 to +10.7 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 02
% OUT-OF-LIMITS = 03
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% PEM 02
% WINDII 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 06
TRANSPONDER FREQUENCIES:
Transponder A: 98/033 - 2287.496961
Transponder B: 98/035 - 2287.499050
GROUND SYSTEM OPERATION
Three (3) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/036 show the observatory data loss to be
20 hours, 41 minutes, 51.540 seconds (an increase of 3.2926 seconds since
last reported). This is a 0.0371 percent data loss which equals a 99.9629
percent data capture for the mission.
The increase of 3.296 seconds occurred over a three-day period as follows:
7 00.960 seconds on Day 030 due to hits in the data,
7 01.792 seconds on Day 034 due to hits in the data, and
7 00.544 seconds on Day 036 due to hits in the data.
No testing of the new CMS software release (R7B1V2) occurred this week.
FOT continued testing IP return-link telemetry this week. Similar to last
week, the FOT utilized a RparallelS AP configuration to monitor the IP
data, which comes into MSOCC on the MS440 (I-channel) and MS442 (Q-channel)
lines. Thus far, the FOT has verified successful I-channel telemetry, but
still has been unable to obtain Q-channel OBC dump data. MSOCC engineering
is aware of the Q-channel problem and is working with NASCOM to rectify
this situation.
Operational processing of level-0 telemetry analysis files continues in
parallel on both the UMPGAL (new MPG Alpha machine) and CDHF computer
systems. No level-0 file transfer failures (between CDHF and UMPGAL)
occurred this week.
SUSIM operations has submitted a request for additional SOLSTICE offsets to
occur on GMT Days 038-043.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued. The
coding is nearly complete.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
FOT attended the MDM/IP Transition Testing Status meeting on 02 February
1998. The purpose of the meeting was to provide the test directors with
feedback on how the individual POCC IP testing has been going. It appears
that most POCCs are unable to get consistently error-free telemetry using
the IP data stream. Only one POCC has attempted commanding using the IP
system. The projected deadline for complete transition to the IP network
is scheduled for mid-April 1998, but it appears that this deadline will
most likely slip to a later date.
FOT met with the MSOCC AP Software programmer (E.Z. Block) on 06 February
1998 for further discussions on adjusting the ground system software to
handle reduced-sized ATC load comparisons. FOT provided additional
information about the process including ATC maps of the regular-sized and
reduced-sized loads.
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 25 February
1998.
UARS Flight Operations plans to attend the Program Experience portion of
the Space Power Workshop at JPL in April 1998. UARS FOT plans on
presenting an overview of the UARS batteries and their performance.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
will probably be required in 1999.
FACILITIES
no change
PERSONNEL CHANGES
Joe Castiglione transferred to UARS from the GGS Control Center on 02
February 1998 to fill an Operations Support Engineer position.
Mike Kokkalakis, Off-Line Engineer, submitted his resignation effective 20
February 1998. Mike accepted a position as Attitude Determination and
Control Engineer with World Space Corporation.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 06 February 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
34929 032/0529 98-038 TMON Cleared TMONs
34941 032/2330 98-039 PEM Power OFF AXIS
34942 033/0147 98-040 OBC Clk Adj fr 20.6 to 41.2
34955 033/2231 98-041 PEM PEM ZEPS Powered Off
34956 034/0009 98-042 WINDII Power ON
34961 034/0808 98-043 WINDII Power OFF
34964 034/1324 98-044 WINDII Power ON
35004 037/0555 98-045 OBC/CLK Clk Adj fr 41.2 to 20.6
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 06 February 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
ACHTRV2 Y-LO ACRIM Heater Voltage 2 34935-41
ACHTRV2 R-LO ACRIM Heater Voltage 2 34943-35016
QTZEPSIF2 Y-LO ZEPS Interface Temperature 2 34948-50
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WIFWPOSN WINDII Filter Wheel Position 34958-62
WISRCBBS WINDII Broadband Source Operating 34963
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 06 February 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
34942 033/01:52 (02 Feb 98) from 20.6 to 41.2 msec/day
35004 037/06:00 (06 Feb 98) from 41.2 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 06 February 1998
31 Jan 1998 - GMT Day 031 Beta = 60.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/ +0.0 18.6 0.000 0.0
2 +0.0/ -78.4 +5.15/+3.55 +25.2/-15.2 26.4 1.027 20.5
3 +56.0/ -11.2 +3.68/+2.29 +24.0/-15.2 26.2 1.032 19.5
06 Feb 1998 - GMT Day 037 Beta = 72.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.6 0.000 0.0
2 +22.4/ -11.2 +3.87/+2.29 +19.2/-16.4 26.6 1.038 17.0
3 +33.6/ -22.4 +2.60/+1.34 +18.8/-16.0 26.6 1.034 16.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
(no change)
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 06 February 1998
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-001 FOT 34924 031/2141 14170 Generic Late Acquisition
98-005 FOT 34942 033/0147 N/A TAC Crashed
98-006 FOT 34962 034/0949 N/A TAC Crashed
98-007 FOT 34972 035/0156 N/A TAC Crashed
98-008 FOT 34988 036/0327 14170 Generic Late Acquisition
98-008 FOT 34992 036/0958 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period